OpenStudy (ale):

what is the solution of the system of equation y=-3x+7 y=2x-8

OpenStudy (anonymous):

what have you got so far?

OpenStudy (anonymous):

Solve the equations simultaneously. Since we have y=-3x+7 and y=2x-8, then -3x+7=2x-8 collecting similar terms together we get 7+8=2x+3x 15=5x x=3. Now substituting x into one of the equations (anyone would do) y=2x-8, where x=3 y=2(3)-8 y=-2. Hence the solutions are x=3 and y=-2